Question 4): If f(x) is differentiable at a, where a > 0, evaluate the following limit in terms of f(a): lim ) – f(a)
Answer #1


* Given a fcw differentiable at &(a) = constant Ilm ta f(n) - fca) f(a) f(a) a - x = a - Ta - Ta = 2 - Indeteminant form Vava